OpenStudy (anonymous):

UMMMM?! HELP:P Write the quadratic equation in factored form. Be sure to write the entire equation. x2 + x - 12 = 0

OpenStudy (anonymous):

x2+x-12=0 x2+4x-3x-12=0 x(x+4)-3(x+4)=0 (x+4)(x-3)=0

OpenStudy (anonymous):

oooooo, i forgot i had to put a = 0, thanks:)